OpenStudy (anonymous):

the lateral area of a right circular cone is 20pi cm^2 and the total area is 36pi cm^2 find the volume of the cone

OpenStudy (cwrw238):

the lateral area of a cone = pi * r * sqrt(h^2 + r^2) area of the base = pi*r^2 and the volume = (1/3)*pi*r^2*h

OpenStudy (cwrw238):

Hint: area of base = 36pi - 20pi = 16pi so pi r^2 (area of base) = 16pi solve for r then plug this value of r into the formula for the lateral area to find the measure of h

OpenStudy (anonymous):

thanks